OpenStudy (anonymous):

A ball is dropped from a height of 20 meters. At what height does the ball have a velocity of 10 meters/second? 5.1 meters 9.8 meters 25 meters 15 meters Done

OpenStudy (anonymous):

E. stripper

OpenStudy (anonymous):

wrong. I failed because of you

OpenStudy (anonymous):

Use \[\color{green}{v_f^2 = v_i^2 + 2gd}\] and solve for d.

OpenStudy (matt101):

You can use the following equation to find your answer: \[v_f^2=v_i^2+2ad\] The final velocity, v(f), is 10 m/s. The initial velocity, v(i), is 0 m/s (the ball starts from rest). The acceleration, a, is that due to gravity, 9.8 m/s^2. The distance, d, is what we're trying to find. So plugging in these numbers gives you: \[10^2=0^2+2(9.8)d \implies d=5.1\] This means the ball travelled 5.1 m down before reaching a speed of 10 m/s. However, this means the ball is still 20-5.1=14.9 m above the ground! The closest answer of your options given is the last one, 15 m, so I'd go with that next time. I'm sorry you didn't get this question right when it mattered, but hopefully you've learned something from these explanations and will do better on your next test!